Bond Valuation

Nesmith Corporation's outstanding bonds have a $1,000 par value, an 8% semiannual coupon, 14 years to maturity, and an 11% YTM. What is the bond's price? (Without using a calculator or excel)
Answer #1

Value of Bond = Interest X PVIFA(r%,n) + Redemption value x PVIF(r%,n)
Interest = 1000 x 8% x 1/2 = 40$
n = no of interest payment = 14 x2 = 28
r = YTM = 11%/2 = 5.5%
Thus value of Bond
= 40xPVIFA(5.5%,28) + 1000 x PVIF(5.5%,28)
= 40 x 14.1214 + 1000 x 0.2233218
= 564.8569 + 223.3218
= 788.1787 $

Note : PVIFA(5.5%,28) and PVIF(5.5%,28) can be found from present value interest factor tables
alternatively it can be found as
PVIF(5.5%,28) = 1/(1+5.5%)^28
= 1/(1.055)^28
=0.2233218

PVIFA(5.5%,28) = [1-(1/(1.055)^28 / 0.055]
= [1-0.2233218 / 0.055]
= 0.776678/0.055
=14.1214